OpenStudy (anonymous):

Use transformation and the zeros of the quadratic function f(x) = (x+4)(x-2) to determine the zeros of each of the following functions.

OpenStudy (anonymous):

a) y=3f(x) and c) y=f(2x)

OpenStudy (campbell_st):

(a) 3f(x) means 3(x+4)(x -2) to find the zeros its just a case of solving 3(x+4)(x-2)=0 then you solve the smaller equations x + 4 = 0 and x - 2 = 0 (b) for the transformation replace x with 2x then f(2x) = (2x + 4)(2x -2) so y = (2x +4)(2x -2) for the zeros then solve (2x +4)(2x -2) = 0 solve the lesser equations 2x + 4 = 0 and 2x -2 = 0

OpenStudy (anonymous):

ohhh. thank you!